OpenStudy (anonymous):

Systems of Linear Equations. Solve by substitution. Solve one of the equations in terms of one of the variables. Substitute this in the other equation. Solve for the variable and replace your answer in the first equation. Check for both in other equation. y=4x+4 y=x-2

OpenStudy (anonymous):

4x + 4 = x - 2 4x - x = -4 -2 3x = -6 x = -6 / 3 x = -2 y = 4(-2) + 4 y = -8 + 4 y = -4

OpenStudy (anonymous):

I do not understand. Can you explain this to me?

OpenStudy (anonymous):

if y is equal to 4x + 4 you simply take the first equation and make it equal to the second equation, then solve for x. Makes sense?

OpenStudy (anonymous):

Yes however I do not understand how you get the 2nd term above.

OpenStudy (anonymous):

how does the x turn into -4?

OpenStudy (callisto):

y=4x+4 -(1) y=x-2 -(2) Sub (1) into (2) 4x + 4 = x - 2 Got it so far?

OpenStudy (anonymous):

x doesnt turn into -4. Once we found x which is -2. we simply go back to the 1st or 2nd equation in the question, and put the value of x in to get the y value of -4

OpenStudy (anonymous):

I understand that part. I guess I'm just confused as to how we found x to = -2. I will review again

OpenStudy (callisto):

4x + 4 = x - 2 Subtract 4 from both sides, you'll get 4x +4 - 4 = x - 2 - 4 4x = x - 6 Subtract x from both sides, you'll get 4x - x = x - 6 - x 3x = -6 Divide both sides by the coefficient of x (that is 3), you'll get 3x / 3 = -6/3 x = -2 Hope you understand it

OpenStudy (anonymous):

I'm starting to understand.... Still reviewing

OpenStudy (anonymous):

I GOT IT. THANKS!!!! I understand. I got confused as to what I was suppose to do.

OpenStudy (anonymous):

good work :)

OpenStudy (callisto):

Glad to hear that you got it :)